OpenStudy (anonymous):

hi<< does cos3x+sin3x=1?? or not?

OpenStudy (anonymous):

square on both sides (cos3x + sin3x)^2 = 1 (cos3x^2 + sin3x^2) + 2 cos3x*sin3x = 1 1 + 2 cos3x*sin3x = 1 2 cos3x*sin3x = 0 so its true only when x=0